# Online Variational Sequential Monte Carlo

### v1

@@ A.1.1. Notation
We now reintroduce the SSM $(X_{t},Y_{t})_{t\in\mathbb{N}}$, specified in Section 2 in the main paper, in a somewhat more rigorous way. More specifically, $(X_{t},Y_{t})_{t\in\mathbb{N}}$ is defined as a bivariate Markov chain evolving on $(\mathsf{X}\times\mathsf{Y},\mathcal{X}\varotimes\mathcal{Y})$ according to the Markov transition kernel
@@
(12) $M_{\theta}:\mathsf{X}\times\mathcal{X}$ $\ni(x,A)\mapsto\int\mathbbm{1}_{A}(x^{\prime})\,m_{\theta}(x,x^{\prime})\,\mu(dx^{\prime}),$
(13) $G_{\theta}:\mathsf{X}\times\mathcal{Y}$ $\ni(x,B)\mapsto\int\mathbbm{1}_{B}(y)\,g_{\theta}(x,y)\,\eta(dy),$
with $m_{\theta}:\mathsf{X}\times\mathsf{X}\to\mathbb{R}_{+}$ and $g_{\theta}:\mathsf{X}\times\mathsf{Y}\to\mathbb{R}_{+}$ being the state and emission transition densities and $\mu\in\mathsf{M}(\mathcal{X})$ and $\eta\in\mathsf{M}(\mathcal{Y})$ reference measures. (Here we have slightly modified the notation in Section 2 of the main paper, by using the short-hand notation $m_{\theta}(x,x^{\prime})=m_{\theta}(x^{\prime}\mid x)$ and $g_{\theta}(x,y)=g_{\theta}(y\mid x)$.) The chain is initialized according to $\chi\varotimes G_{\theta}:\mathcal{X}\varotimes\mathcal{Y}\ni A\mapsto\int_{A}\chi(dx)\,G_{\theta}(x,dy)$, where $\chi$ is some probability measure on $(\mathsf{X},\mathcal{X})$ having density $m_{0}(x)$ with respect to $\mu$. As specified in Section 2, only the process $(Y_{t})_{t\in\mathbb{N}}$ is observed, whereas the state process $(X_{t})_{t\in\mathbb{N}}$ is unobserved and hence referred to as latent or hidden. Moreover, $\theta$ is a parameter belonging to some vector space $\mathsf{\Theta}$ and governing the dynamics of the model.
@@ A.1.2. Algorithm
Algorithm 3 displays the procedure studied in our analysis. In this slightly modified version of Algorithm 2 in the main paper, the particle system will be represented by the resampled particles, denoted here as $(\tilde{\xi}_{t}^{i})_{i=1}^{N}$, $t\geq 0$. In order to avoid introducing further notation, the resampled particles are conventionally all initialized at time $-1$ by some arbitrary value $u\in\mathsf{X}$ such that $\{f_{\theta}(u,y_{0},\varepsilon_{0}^{i})\}_{i=1}^{N}$, corresponding to $(\xi_{0}^{i})_{i=1}^{N}$, are i.i.d. according to some initial proposal distribution $R_{\theta}((u,y_{0}),\cdot)$ depending on $u$, where $(\varepsilon_{0}^{i})_{i=1}^{N}\sim\nu^{\varotimes N}$. As we will see, the cloud of resampled particles will be included in the state-dependent Markov chain $(Z_{t})_{t\in\mathbb{N}}$ governing the perturbations of the stochastic-approximation scheme under consideration; the initialization according to the constant $u$ is described in (15).
In Algorithm 3, we operate with two samples of mutated particles, one used for the propagation of the particle cloud and one used for approximation of the gradient, in the sense that the noise variables $(\varepsilon_{t+1}^{i})_{i=1}^{N}$ generated on Line 8 are not used to propagate the particles. Consequently, $f_{\theta_{t}}(\tilde{\xi}_{t}^{i},y_{t+1},\varepsilon_{t+1}^{i})$ is generally different from $\xi_{t+1}^{i}$. Importantly, this decoupling makes the Markov transition kernel $T_{\theta}$ non-collapsed (free from Dirac components), which, as we will see, facilitates significantly the theoretical analysis of the algorithm.
@@
12:return $(\tilde{\xi}_{t}^{i})_{i=1}^{N}$, $\theta_{t+1}$
@@ A.2.1. Construction of $(Z_{t})_{t\in\mathbb{N}}$
We first provide a more detailed statement of Assumption 4.1 in the main paper, which assumes that the law of the data is governed by an unspecified SSM.
@@
As explained in Section 4 in the main paper, the stochastic process $(Z_{t})_{t\in\mathbb{N}}$, evolving on the product space $(\mathsf{Z},\mathcal{Z})\vcentcolon=(\mathsf{X}\times\mathsf{Y}\times\mathsf{X}^{N}\times\mathsf{E}^{N},\mathcal{X}\varotimes\mathcal{Y}\varotimes\mathcal{X}^{\varotimes N}\varotimes\mathcal{E}^{\varotimes N})$, includes the data-generating SSM well as the random variables generated by Algorithm 3, i.e., for $t\in\mathbb{N}$, $Z_{t}\coloneqq(X_{t},Y_{t},\tilde{\xi}_{t-1}^{1:N},\varepsilon_{t}^{1:N})$. Let $\mathbb{P}$ and $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be the law and natural filtration of $(Z_{t})_{t\in\mathbb{N}}$; then, as described in Section 4, $(Z_{t})_{t\in\mathbb{N}}$ is a state-dependent Markov chain with transition kernel $T_{\theta}$, in the sense that for any bounded measurable function $h$ on $\mathsf{Z}$, $\mathbb{P}$-a.s., $\mathbb{E}[h(Z_{t+1})\mid\mathcal{F}_{t}]=T_{\theta_{t}}h(Z_{t})$. The kernel $T_{\theta}$ is given by, with $z_{t}=(x_{t},y_{t},\tilde{x}_{t-1}^{1:N},v_{t}^{1:N})$,
(14) $T_{\theta}(z_{t},dz_{t+1})\vcentcolon=\bar{M}(x_{t},dx_{t+1})\,\bar{G}(x_{t+1},dy_{t+1})\int_{x_{t}^{1:N}}\prod_{k=1}^{N}r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t})\\ \times\prod_{j=1}^{N}\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)\,\nu^{\varotimes N}(dv_{t+1}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N}),$
@@
(15) $\tau_{0}(dz_{0})\vcentcolon=\chi(dx_{0})\,\bar{G}(x_{0},dy_{0})\,\delta_{u}^{\varotimes N}(d\tilde{x}_{-1}^{1:N})\,\nu^{\varotimes N}(dv_{0}^{1:N}),$
@@
We now rewrite and prove Proposition 4.3 in the main paper.
@@
(16) $T_{\theta}(z_{t},A)$ =∫⋯∫$\idotsint\mathbbm{1}_{A}(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})= ∫⋯$
(17) ×$\qquad\times\int_{x_{t}^{1:N}}\prod_{k=1}^{N}r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t})\prod_{j=1}^{N}\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)$
(18) ×$\hskip 227.62204pt\times\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N})$
(19) $\geq\epsilon^{7N+1}\int\cdots\int\mathbbm{1}_{A}(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})$
(20) ×$\qquad\times\int_{x_{t}^{1:N}}\prod_{j=1}^{N}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N}).$
@@
(21) $\nu_{1}(dz)=\epsilon^{7N+1}\bar{g}(x,y)\,\mu(dx)\,\eta(dy)\,\int_{x^{1:N}}\prod_{j=1}^{N}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}(d\tilde{x}^{j})\right)\,\mu^{\varotimes N}(dx^{1:N})\,\nu^{\varotimes N}(dv^{1:N}).$
Then, by (Meyn and Tweedie,, 2009, Theorem 16.0.2(v)) or (Douc et al.,, 2018, Theorem 15.3.1(iii)) it follows that $(Z_{t}^{\theta})_{t\geq 0}$ is geometrically uniformly ergodic. Then the Dobrushin coefficient of $T_{\theta}$ is strictly less than one for all $\theta\in\mathsf{\Theta}$, which implies that $T_{\theta}$ admits a unique stationary distribution $\tau_{\theta}$ (see, e.g., Douc et al.,, 2018, Theorems 18.2.4–5). ∎
@@ A.2.2. Stochastic approximation update and mean field
(22) $H_{\theta}(z)\coloneqq\nabla\ln\left(\sum_{i=1}^{N}w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)\right)=\frac{\sum_{i=1}^{N}\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)}{\sum_{i=1}^{N}w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)},$
@@
(23) $\max\{\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))\rVert,\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)\rVert,\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\}\leq\kappa,$
(24) $\max\{\left\lvert m_{\theta}(x,x^{\prime})-m_{\theta^{\prime}}(x,x^{\prime})\right\rvert,\left\lvert m_{\theta}(x,f_{\theta}(x,y,v))-m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\right\rvert,$
(25) $\hskip 85.35826pt\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))-\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert,$
(26) $\max\{\left\lvert g_{\theta}(x,y)-g_{\theta^{\prime}}(x,y)\right\rvert,\left\lvert g_{\theta}(f_{\theta}(x,y,v),y)-g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\right\rvert,$
(27) $\hskip 85.35826pt\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)-\nabla g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert,$
(28) $\max\{\left\lvert r_{\theta}(x,x^{\prime},y)-r_{\theta^{\prime}}(x,x^{\prime},y)\right\rvert,\left\lvert r_{\theta}(x,f_{\theta}(x,y,v),y)-r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\right\rvert,$
(29) $\hskip 85.35826pt\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert.$
@@
    (30) $H_{\theta}(z)-h(\theta)=\tilde{H}_{\theta}(z)-T_{\theta}\tilde{H}_{\theta}(z),$
@@
    (31) $\max\{\lVert H_{\theta}(z)\rVert,\lVert\tilde{H}_{\theta}(z)\rVert,\lVert T_{\theta}\tilde{H}_{\theta}(z)\rVert\}\leq\tilde{\alpha},$
    (32) $\max\{\lVert H_{\theta}(z)-H_{\theta^{\prime}}(z)\rVert,\lVert T_{\theta}\tilde{H}_{\theta}(z)-T_{\theta^{\prime}}\tilde{H}_{\theta^{\prime}}(z)\rVert,\lVert h(\theta)-h(\theta^{\prime})\rVert\}\leq\tilde{\alpha}\lVert\theta-\theta^{\prime}\rVert.$

### v3

@@ A.1.1. Notation
We now reintroduce the SSM $(X_{t},Y_{t})_{t\in\mathbb{N}}$, specified in Section 2, in a somewhat more rigorous way. More specifically, $(X_{t},Y_{t})_{t\in\mathbb{N}}$ is defined as a bivariate Markov chain evolving on $(\mathsf{X}\times\mathsf{Y},\mathcal{X}\varotimes\mathcal{Y})$ according to the Markov transition kernel
@@
$M_{\theta}:\mathsf{X}\times\mathcal{X}$ $\ni(x,A)\mapsto\int\mathbbm{1}_{A}(x^{\prime})\,m_{\theta}(x,x^{\prime})\,\mu(dx^{\prime}),$ (10)
$G_{\theta}:\mathsf{X}\times\mathcal{Y}$ $\ni(x,B)\mapsto\int\mathbbm{1}_{B}(y)\,g_{\theta}(x,y)\,\eta(dy),$ (11)
with $m_{\theta}:\mathsf{X}\times\mathsf{X}\to\mathbb{R}_{+}$ and $g_{\theta}:\mathsf{X}\times\mathsf{Y}\to\mathbb{R}_{+}$ being the state and emission transition densities and $\mu\in\mathsf{M}(\mathcal{X})$ and $\eta\in\mathsf{M}(\mathcal{Y})$ reference measures. (Here we have slightly modified the notation in Section 2, by using the short-hand notation $m_{\theta}(x,x^{\prime})=m_{\theta}(x^{\prime}\mid x)$ and $g_{\theta}(x,y)=g_{\theta}(y\mid x)$.) The chain is initialized according to $\chi\varotimes G_{\theta}:\mathcal{X}\varotimes\mathcal{Y}\ni A\mapsto\int_{A}\chi(dx)\,G_{\theta}(x,dy)$, where $\chi$ is some probability measure on $(\mathsf{X},\mathcal{X})$ having density $m_{0}(x)$ with respect to $\mu$. As specified in Section 2, only the process $(Y_{t})_{t\in\mathbb{N}}$ is observed, whereas the state process $(X_{t})_{t\in\mathbb{N}}$ is unobserved and hence referred to as latent or hidden. Moreover, $\theta$ is a parameter belonging to some vector space $\mathsf{\Theta}$ and governing the dynamics of the model.
@@ A.1.2. Algorithm
Algorithm 3 displays the procedure studied in our analysis. In this slightly modified version of Algorithm 2, the particle system will be represented by the resampled particles, denoted here as $(\tilde{\xi}_{t}^{i})_{i=1}^{N}$, $t\geq 0$. In order to avoid introducing further notation, the resampled particles are conventionally all initialized at time $-1$ by some arbitrary value $u\in\mathsf{X}$ such that $\{f_{\theta}(u,y_{0},\varepsilon_{0}^{i})\}_{i=1}^{N}$, corresponding to $(\xi_{0}^{i})_{i=1}^{N}$, are i.i.d. according to some initial proposal distribution $R_{\theta}((u,y_{0}),\cdot)$ depending on $u$, where $(\varepsilon_{0}^{i})_{i=1}^{N}\sim\nu^{\varotimes N}$. As we will see, the cloud of resampled particles will be included in the state-dependent Markov chain $(Z_{t})_{t\in\mathbb{N}}$ governing the perturbations of the stochastic-approximation scheme under consideration; the initialization according to the constant $u$ is described in (13).
In Algorithm 3, we operate with two samples of mutated particles, one used for the propagation of the particle cloud and one used for approximation of the gradient, in the sense that the noise variables $(\varepsilon_{t+1}^{i})_{i=1}^{N}$ generated on Line 9 are not used to propagate the particles. Consequently, $f_{\theta_{t}}(\tilde{\xi}_{t}^{i},y_{t+1},\varepsilon_{t+1}^{i})$ is generally different from $\xi_{t+1}^{i}$. Importantly, this decoupling makes the Markov transition kernel $T_{\theta}$ non-collapsed (free from Dirac components), which, as we will see, facilitates significantly the theoretical analysis of the algorithm.
@@
12:  return $(\tilde{\xi}_{t}^{i})_{i=1}^{N}$, $\theta_{t+1}$.
@@ A.2.1. Construction of $(Z_{t})_{t\in\mathbb{N}}$
We first provide a more detailed statement of Assumption 4.1, which assumes that the law of the data is governed by an unspecified SSM.
@@
As explained in Section 4, the stochastic process $(Z_{t})_{t\in\mathbb{N}}$, evolving on the product space $(\mathsf{Z},\mathcal{Z})\vcentcolon=(\mathsf{X}\times\mathsf{Y}\times\mathsf{X}^{N}\times\mathsf{E}^{N},\mathcal{X}\varotimes\mathcal{Y}\varotimes\mathcal{X}^{\varotimes N}\varotimes\mathcal{E}^{\varotimes N})$, includes the data-generating SSM well as the random variables generated by Algorithm 3, i.e., for $t\in\mathbb{N}$, $Z_{t}\coloneqq(X_{t},Y_{t},\tilde{\xi}_{t-1}^{1:N},\varepsilon_{t}^{1:N})$. Let $\mathbb{P}$ and $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be the law and natural filtration of $(Z_{t})_{t\in\mathbb{N}}$; then, as described in Section 4, $(Z_{t})_{t\in\mathbb{N}}$ is a state-dependent Markov chain with transition kernel $T_{\theta}$, in the sense that for any bounded measurable function $h$ on $\mathsf{Z}$, $\mathbb{P}$-a.s., $\mathbb{E}[h(Z_{t+1})\mid\mathcal{F}_{t}]=T_{\theta_{t}}h(Z_{t})$. The kernel $T_{\theta}$ is given by, with $z_{t}=(x_{t},y_{t},\tilde{x}_{t-1}^{1:N},v_{t}^{1:N})$,
$T_{\theta}(z_{t},dz_{t+1})\vcentcolon=\bar{M}(x_{t},dx_{t+1})\,\bar{G}(x_{t+1},dy_{t+1})\int_{x_{t}^{1:N}}\prod_{k=1}^{N}r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t})\\ \times\prod_{j=1}^{N}\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)\,\nu^{\varotimes N}(dv_{t+1}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N}),$ (12)
@@
$\tau_{0}(dz_{0})\vcentcolon=\chi(dx_{0})\,\bar{G}(x_{0},dy_{0})\,\delta_{u}^{\varotimes N}(d\tilde{x}_{-1}^{1:N})\,\nu^{\varotimes N}(dv_{0}^{1:N}),$ (13)
@@
We now rewrite and prove Proposition 4.3.
@@
$T_{\theta}(z_{t},A)$ =∫⋯∫$\idotsint\mathbbm{1}_{A}(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})= ∫⋯$ (14)
×$\qquad\times\int_{x_{t}^{1:N}}\prod_{k=1}^{N}r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t})\prod_{j=1}^{N}\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)$ (15)
×$\hskip 227.62204pt\times\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N})$ (16)
$\geq\epsilon^{7N+1}\int\cdots\int\mathbbm{1}_{A}(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})$ (17)
×$\qquad\times\int_{x_{t}^{1:N}}\prod_{j=1}^{N}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i}}(d\tilde{x}_{t}^{j})\right)\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N}).$ (18)
@@
$\nu_{1}(dz)=\epsilon^{7N+1}\bar{g}(x,y)\,\mu(dx)\,\eta(dy)\,\int_{x^{1:N}}\prod_{j=1}^{N}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}(d\tilde{x}^{j})\right)\,\mu^{\varotimes N}(dx^{1:N})\,\nu^{\varotimes N}(dv^{1:N}).$ (19)
Then, by Meyn and Tweedie, (2009, Theorem 16.0.2(v)) or Douc et al., (2018, Theorem 15.3.1(iii)) it follows that $(Z_{t}^{\theta})_{t\geq 0}$ is geometrically uniformly ergodic. Then the Dobrushin coefficient of $T_{\theta}$ is strictly less than one for all $\theta\in\mathsf{\Theta}$, which implies that $T_{\theta}$ admits a unique stationary distribution $\tau_{\theta}$ (see, e.g., Douc et al.,, 2018, Theorems 18.2.4–5). ∎
@@ A.2.2. Stochastic approximation update and mean field
$H_{\theta}(z)\coloneqq\nabla\ln\left(\sum_{i=1}^{N}w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)\right)=\frac{\sum_{i=1}^{N}\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)}{\sum_{i=1}^{N}w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)},$ (20)
@@
$\max\{\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))\rVert,\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)\rVert,\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\}\leq\kappa,$ (21)
$\max\{\left\lvert m_{\theta}(x,x^{\prime})-m_{\theta^{\prime}}(x,x^{\prime})\right\rvert,\left\lvert m_{\theta}(x,f_{\theta}(x,y,v))-m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\right\rvert,$ (22)
$\hskip 85.35826pt\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))-\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert,$ (23)
$\max\{\left\lvert g_{\theta}(x,y)-g_{\theta^{\prime}}(x,y)\right\rvert,\left\lvert g_{\theta}(f_{\theta}(x,y,v),y)-g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\right\rvert,$ (24)
$\hskip 85.35826pt\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)-\nabla g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert,$ (25)
$\max\{\left\lvert r_{\theta}(x,x^{\prime},y)-r_{\theta^{\prime}}(x,x^{\prime},y)\right\rvert,\left\lvert r_{\theta}(x,f_{\theta}(x,y,v),y)-r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\right\rvert,$ (26)
$\hskip 85.35826pt\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert\}\leq\kappa\lVert\theta-\theta^{\prime}\rVert.$ (27)
@@
    $H_{\theta}(z)-h(\theta)=\tilde{H}_{\theta}(z)-T_{\theta}\tilde{H}_{\theta}(z),$ (28)
@@
    $\max\{\lVert H_{\theta}(z)\rVert,\lVert\tilde{H}_{\theta}(z)\rVert,\lVert T_{\theta}\tilde{H}_{\theta}(z)\rVert\}\leq\tilde{\alpha},$ (29)
    $\max\{\lVert H_{\theta}(z)-H_{\theta^{\prime}}(z)\rVert,\lVert T_{\theta}\tilde{H}_{\theta}(z)-T_{\theta^{\prime}}\tilde{H}_{\theta^{\prime}}(z)\rVert,\lVert h(\theta)-h(\theta^{\prime})\rVert\}\leq\tilde{\alpha}\lVert\theta-\theta^{\prime}\rVert.$ (30)